SO_{2} SO_{2}

SO_{2}

X̃_{i} = X_{i}+ bϵ_{i},

X_{i} SO_{2} b =

1.03σ̂n^{−1/3} ϵ_{i}, i = 1, …, n

(−\sqrt{5}, \sqrt{5})

σ̂ b

θ̃_{i} = θ_{i}+ dϵ_{i},

θ_{i} ϵ_{i}, i = 1, …, n μ = 0 κ = 1 d =

n^{−1/5} ?? ?? [0, 2π)

\frac{π}{2} π \frac{3π}{2}

??

R̄

?? ?? ?? ?? ?? R̄

?? de_{r}osa.epsDiagramaderosadelasmediashorariasdeladireccindeviento. ?? ?? ?? ?? ??

????

??

?? ?? ?? ?? ?? ?? ?? ?? ?? ?? ?? no_{p}aramT heta1.epsEstimacinnoparamtricatiponcleodelafuncindedensidadcirculardelasmediashorariasdeladireccindevientocorrespondientesalprimerperiodode2010.P armetrodesuavizadoobtenidomedianteplug − in(violeta), mediantevalidacincruzadadadaporlafuncindeprdidaL2(azul)yporlafuncindeprdidaKullback − Leibler(verde). no_{p}aramT heta2.epsEstimacinnoparamtricatiponcleodelafuncindedensidadcirculardelasmediashorariasdeladireccindevientocorrespondientesalsegundoperiodode2010.P armetrodesuavizadoobtenidomedianteplug − in(violeta), mediantevalidacincruzadadadaporlafuncindeprdidaL2(azul)yporlafuncindeprdidaKullback − Leibler(verde). no_{p}aramT heta3.epsEstimacinnoparamtricatiponcleodelafuncindedensidadcirculardelasmediashorariasdeladireccindevientocorrespondientesaltercerperiodode2010.P armetrodesuavizadoobtenidomedianteplug − in(violeta), mediantevalidacincruzadadadaporlafuncindeprdidaL2(azul)yporlafuncindeprdidaKullback − Leibler(verde). no_{p}aramT heta4.epsEstimacinnoparamtricatiponcleodelafuncindedensidadcirculardelasmediashorariasdeladireccindevientocorrespondientesalcuartoperiodode2010.P armetrodesuavizadoobtenidomedianteplug − in(violeta), mediantevalidacincruzadadadaporlafuncindeprdidaL2(azul)yporlafuncindeprdidaKullback − Leibler(verde).

?? Θ Φ

ρ_{T} =

E[(Θ_{1}−Θ_{2})(Φ_{1}−Φ_{2})] ; {E[^{2}(Θ_{1}−Θ_{2})]E[^{2}(Φ_{1}−Φ_{2})]}^{1/2}

(Θ_{1}, Φ_{1}) (Θ_{2}, Φ_{2})

(Θ, Φ) ρ_{T}

ρ̂_{T} =

\frac{∑_{1≤i<j≤n}(θ_{i}−θ_{j})(ϕ_{i}−ϕ_{j})}{[∑_{1≤i<j≤n}^{2}(θ_{i}−θ_{j}) ∑_{1≤i<j≤n}^{2}(ϕ_{i}−ϕ_{j})]^{1/2}}

ϕ_{i} = θ_{i+k}, i = 1, …, n−

k (n−

k) ϕ_{1}, θ_{1}), (ϕ_{2}, θ_{2}), …, (ϕ_{n−k}, θ_{n−k})

ρ̂_{T}^{k}

ρ̂_{T}^{k} =

\frac{∑_{1≤i<j≤n−k}(θ_{i}−θ_{j})(ϕ_{i}−ϕ_{j})}{[∑_{1≤i<j≤n−k}^{2}(θ_{i}−θ_{j}) ∑_{1≤i<j≤n−k}^{2}(ϕ_{i}−ϕ_{j})]^{1/2}}

ρ̂^{k}_{T} = \frac{∑_{1≤i<j≤n−k}(θ_{i} − θ_{j})(θ_{i+k} − θ_{j+k})}{[∑_{1≤i<j≤n}^{2}(θ_{i} − θ_{j}) ∑_{1≤i<j≤n−k}^{2}(θ_{i+k} − θ_{j+k})]^{1/2}}

(1)

ρ̂_{T}^{k}

?? ?? ?? ??

(?? α ω κ ?? ω ∈ [−1, 1] −πα < π ?? ?? ?? ??) / (\overline{n ∑_{i=1}^{n} d(θ_{i},θ̂_{i})^{2}= \frac{1}{n} ∑_{i=1}^{n}(π−|π−|θ_{i}−θ̂_{i}||)^{2}})

θ_{i}, i = 1, …, n

θ̂_{i}, i = 1, …, n

d

?? ?? ?? ?? π

αωκ

ts_{k}T heta1.epsModelodeMbiusdeseriesdetiempodelasmediashorariasdeladireccindevientocorrespondientesalprimerperiodode2010. ts_{k}T heta2.epsModelodeMbiusdeseriesdetiempodelasmediashorariasdeladireccindevientocorrespondientesalsegundoperiodode2010. ts_{k}T heta3.epsModelodeMbiusdeseriesdetiempodelasmediashorariasdeladireccindevientocorrespondientesaltercerperiodode2010. ts_{k}T heta4.epsModelodeMbiusdeseriesdetiempodelasmediashorariasdeladireccindevientocorrespondientesalcuartoperiodode2010

2 2 2 2

?? ??

?? ?? ?? ?? ?? ?? ?? ??

κ = 1/h_{s}^{2}

2

?? ?? ?? ?? ??

2 2 2

?? ?? ?? ?? ??

2 2 2

??

2 2 2 2

??

2 2

?? ?? ?? ?? ?? ?? ?? ?? ?? ?? comp2B1.epsEstimacindelafuncinderegresincircular − linealenB1.Eleccindelparmetrodesuavizadomediante_{s}^{2} comp3B1.epsEstimacindelafuncinderegresincircular − linealenB1.Eleccindelparmetrodesuavizadomediante_{s}^{2} comp2B2.epsEstimacindelafuncinderegresincircular − linealenB2.Eleccindelparmetrodesuavizadomediante_{s}^{2} comp3B2.epsEstimacindelafuncinderegresincircular − linealenB2.Eleccindelparmetrodesuavizadomediante_{s}^{2} comp2C9.epsEstimacindelafuncinderegresincircular − linealenC9.Eleccindelparmetrodesuavizadomediante_{s}^{2} comp3C9.epsEstimacindelafuncinderegresincircular − linealenC9.Eleccindelparmetrodesuavizadomediante_{s}^{2} comp2F 2.epsEstimacindelafuncinderegresincircular − linealenF 2.Eleccindelparmetrodesuavizadomediante_{s}^{2} comp3F 2.epsEstimacindelafuncinderegresincircular − linealenF 2.Eleccindelparmetrodesuavizadomediante_{s}^{2} comp2G2.epsEstimacindelafuncinderegresincircular − linealenG2.Eleccindelparmetrodesuavizadomediante_{s}^{2}